OpenStudy (anonymous):

8. Solve for x: 4(x−3)+2x=3(2x−5) A. x = 0 B. x = all real numbers C. x = 6 D. there is no solution

OpenStudy (anonymous):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

you take the number or sign right before the parenthesis and multiply it by everything in the parenthesis. 4(x - 3) + 2x = 3(2x - 5) 4x - 12 + 2x = 6x - 15 -- combine like terms 6x - 12 = 6x - 15 -- add 12 to both sides 6x = 6x - 15 + 12 -- combine like terms 6x = 6x - 3 -- subtract 6x from both sides 0 = -3 -- incorrect There is no solution any questions ?